# Hugo Malard1  Michel Olvera1  Stéphane Lathuiliere1  Slim Essid1 1LTCI, Télécom Paris, Institut Polytechnique de Paris

### v1

@@ 5.2 Sound-Prompted Semantic Segmentation
Sound-Prompted Semantic Segmentation differs from Sound-Prompted Segmentation seen in the previous section as it requires the estimation of class labels corresponding to the mask. In TACO, classifying the audio prompt $V_{I}^{k^{\star}}$ allows us to obtain the label of the segment. We run experiments on the AVSS and ADE20k sound-prompted semantic datasets. As shown in Table 5, our method greatly outperforms the baseline on the two datasets as it can take advantage of both the image and the audio decompositions (+8 and 15 points of mIoU in AVSS and ADE SP Semantic, respectively). On the ADE SP Semantic, which contains a single class per image, the method performs well. However, on the AVSS dataset that contains samples with multiple classes, the performances drop. We attribute this to the class decoding process: since we estimate the class label as the closest one to the value of $V_{I}^{k^{\star}}$, we consider the full segmentation as a single class. Nonetheless, the class-agnostic segmentation masks are accurate as the decomposition encodes all the concepts in a single vector (shown in supplementary materials).
@@ 5.3 Qualitative Analysis
Figure 4 shows images along with their associated $V_{I}^{k^{\star}}$ reshaped as a matrix, the segmentation from TACO, and an illustration of the associated audio. The first row shows a challenging example from the S4 dataset where the sound source is nearly invisible, yet both TACO and $V_{I}^{k^{\star}}$ successfully segment it. While recent works [20] underline how audio-visual segmenter models fail to deal with negative audio (i.e. audio depicting an event not present in the image), the second row illustrates the robustness of our method on that particular problem. As there is no common concept between the audio and image (given that the sound track of this example features music instead of the original audio recording of the scene), the estimated $V_{I}^{k^{\star}}$ contains only noise, and as FC-CLIP does not find the concept, hence the final segmentation is empty, showing the resilience of TACO to those negative audios. Finally, the last row illustrates the case of multiple potential sources in the image. Even though the image features a Guitar, a Djembe, and singers, the model correctly segments only the people as only singing voice is heard on this excerpt.
@@
Figure 4: Qualitative segmentation examples from ADE SP and S4 datasets.
@@
Analysis in the Multi-Source setting. Figure 5 shows a segmentation example from the MS3 dataset as well as the similarity of the associated $V_{I}^{k^{\star}}$ with some words. In the first row, the sound heard is a baby babbling, while the image contains both the baby, a woman, and a sofa. As expected, the closest word to $V_{I}^{k^{\star}}$ “Baby babbling”, however, in the second row (next second of the video) both the women and the baby are emitting sound. As one might have expected that those two sounds were encoded in two different factors, it turns out that they both are encoded on the sounding factor $V_{I}^{k^{\star}}$, as its similarity is maximal with “A baby babbling and a woman talking”, which is coherent with the segmentation. This observation enlightens our previous hypothesis made in section 5.1: the decomposition does not only encode one common concept between the audio and image features, but all the sounding objects, explaining its great performances in the multi-source task.
Figure 5: Multiple source segmentation examples. First row, a baby babbling is correctly segmented. Second row, the sounds of both a baby and a woman are encoded in a single factor showing that our decomposition encodes all sounding objects together.
6 Conclusion
@@
We introduced TACO, a method for training-free sound-prompted segmentation that effectively prompts an open vocabulary image segmentation model such as FC-CLIP with factors extracted through co-factorization of audio-visual deep embedding representations. Our method provides a framework for unsupervised sound-prompted segmentation, enabling the extraction and visualization of shared sounding concepts across audio spectrograms and image regions, which makes it inherently interpretable. TACO significantly outperforms concurrent unsupervised approaches, achieving state-of-the-art results.
Supplementary Material
The supplementary material is organized as follows: the first part presents an analysis of the computations induced by the decomposition method, followed by an explanation of the metrics used and a comparison with other works. The last part consists of qualitative visualization and analysis of segmentation from TACO.
7 Computational overhead
@@
8 Metrics discussion
@@
9 Semantic segmentation analysis
@@
10 Qualitative results of single and multiple sources, sound-prompted segmentation
@@
11 Typical failure cases

### v3

@@ 4.2 Sound-Prompted Semantic Segmentation
Sound-Prompted Semantic Segmentation differs from Sound-Prompted Segmentation seen in the previous section as it requires the estimation of class labels corresponding to the mask. In TACO, classifying the audio prompt $V_{I}^{k^{\star}}$ allows us to obtain the label of the segment. We run experiments on the AVSS and ADE20k SP Semantic datasets. As shown in Table 3, our method greatly outperforms the baseline on the two datasets as it can take advantage of both the image and the audio decompositions (+8 and 15 points of mIoU in AVSS and ADE SP Semantic, respectively).
On the ADE SP Semantic, which contains a single class per image, the method performs well. However, on the AVSS dataset that contains samples with multiple classes, the performances drop. We attribute this to the class decoding process: since we estimate the class label as the closest one to the value of $V_{I}^{k^{\star}}$, we consider the full segmentation as a single class. Nonetheless, the class-agnostic segmentation masks are accurate as the decomposition encodes all the concepts in a single vector (shown in Appendix G).
@@ 4.3 Qualitative Analysis
Figure 5 shows images along with their associated $V_{I}^{k^{\star}}$ reshaped as a matrix, the segmentation from TACO, and an illustration of the associated audio. These examples highlight both the segmentation quality of TACO and the method’s interpretability. Specifically, $U_{I}^{k^{\star}}$ reveals the tokens attended by the decomposition, from which $V_{I}^{k^{\star}}$ serves as the prompt for FC-CLIP. The first row shows a challenging example from the S4 dataset, where the sound source is nearly invisible, yet both TACO and $V_{I}^{k^{\star}}$ successfully segment it. While recent works [21] underline how audio-visual segmenter models fail to deal with negative audio (i.e. audio depicting an event not present in the image), the second row illustrates the robustness of our method on that particular problem. As there is no common concept between the audio and image (given that the sound track of this example features music instead of the original audio recording of the scene), the estimated $V_{I}^{k^{\star}}$ contains only noise, therefore FC-CLIP does not find it in the image, hence the final segmentation is empty, showing the resilience of TACO to those negative audios. Finally, the last row illustrates the case of multiple potential sources in the image. Even though the image features a Guitar, a Djembe, and singers, the model correctly segments only the people as only singing voice is heard on this excerpt.
@@
Figure 4: Qualitative segmentation examples from ADE SP and S4 datasets: TACO performs well even in challenging cases
Figure 5: Multiple source segmentation examples. As the sources change during the video, TACO’s segmentation changes as well
@@
Figure 5 shows a segmentation example from the multi-source MS3 dataset. To gain insight into the model’s behavior, we manually selected words describing sounding entities in the image and examined the similarity $V_{I}^{k^{\star}}$ with them. In the first row, the sound heard is a baby babbling ([Uncaptioned image] in the first column), while the image contains both the baby, a woman, and a sofa. As expected, the closest word to $V_{I}^{k^{\star}}$ is “Baby babbling" (see [Uncaptioned image] in the last column). In the second row, both the women and the baby are making sounds ([Uncaptioned image]+[Uncaptioned image] in the first column). While one might expect these two sounds to be encoded in two different factors, they are both encoded in the same sounding factor $V_{I}^{k^{\star}}$, as its similarity is maximal with “A baby babbling and a woman talking" ([Uncaptioned image]+[Uncaptioned image] in the last column), which is coherent with the segmentation. This observation reinforces our hypothesis made in section 4.1: the decomposition encodes not just a common concept between audio and image features, but all the sounding objects, explaining its strong performance in the multi-source task.
5 Conclusion and limitations
@@
We introduced TACO, a method for training-free sound-prompted visual segmentation effectively incorporating an open vocabulary image segmentation model, namely FC-CLIP, by prompting it with factors extracted through a semantically motivated co-factorization of audio-visual deep embedding representations. Our method provides a framework for unsupervised audio-visual segmentation, enabling the extraction and visualization of sounding concepts shared across audio spectrograms and image regions, making it inherently interpretable. TACO outperforms concurrent approaches, achieving state-of-the-art results. However, although the Sem CO-NMF incurs lower costs compared to pretrained models, it still introduces additional computational overhead during inference, which may pose challenges in resource-constrained settings.
Appendix A Analysis of clamping resilience
To assess the resilience of CLIP representations to clamping negative values, we evaluate FC-CLIP using only the nonnegative portion (obtained via ReLU) of the CLIP text encoding, a variant referred to as A+-FC-CLIP. Table 4 shows the performance degradation across all datasets. Except for the ADE SP dataset ($\sim$3.5 points drop), the decrease never exceeds 2 points. This experiment validates the use of clamping, enabling the application of NMF-based methods so as to benefit from their interpretability.
ADE SP S4 MS3 AVSS ADE SP S.
A-FC-CLIP 45.31 51.52 33.67 11.95 24.95
A+-FC-CLIP 41.89 51.63 31.67 11.82 23.73
Appendix B Analysis of hyper-parameters
Table 6 shows the performance of our method on the S4 dataset using different numbers of components in the decomposition (values of K). Results are stable across multiple values, showing the resilience of the method with respect to the number of factors.
Table 6 shows the performance of TACO (on the S4 dataset) with respect to the hyperparameter $\beta_{p}$. The performance remaining stable across multiple values of the hyperparameter indicates the robustness of the method. TACO-KL shows the performance of TACO when using the KL-divergence instead of the cross entropy to measure dissimilarity between the semantic descriptors. TACO Word/2 and Word/4 respectively correspond to TACO using half and a quarter of the words in the general word bank (randomly selected). Using KL-divergence instead of cross entropy affects significantly the performance of TACO. Interestingly dividing by 2 the word bank does not change much the performance, however, when dividing by four, performance starts to drop a lot.
S4
K mask-IoU $\uparrow$ F-Score $\uparrow$
6 61.74 68.87
8 64.04 71.50
10 63.92 71.54
12 63.78 71.28
Table 5: Variation of the number of factors
S4
Method $\beta_{p}$ mask-IoU $\uparrow$ F-Score $\uparrow$
100 63.45 70.57
TACO 125 64.04 71.50
150 63.90 71.52
TACO-KL 125 58.74 65.69
TACO Words/2 125 62.91 70.16
TACO Words/4 125 55.50 62.10
Table 6: Variation of the design choices
Word Bank S4 MS3 ADE SP
$U_{i}^{k^{\star}}$ General 27.78$\pm$0.06 25.88$\pm$0.91 32.42$\pm$0.74
Specific 28.76$\pm$0.06 26.57$\pm$0.17 27.74$\pm$0.48
TACO Specific 64.73$\pm$0.16 45.15$\pm$0.80 54.05$\pm$1.40
General 64.04$\pm$0.25 43.15$\pm$0.91 51.57$\pm$0.86
Table 7 presents the mask-IoU scores on the three used sound-prompted segmentation datasets using either a general class list or one specific to the dataset (i.e. the list of all the classes of the dataset). Using a specific class list overall improves both TACO results (+1 point on S4, 3.5 points on ADE SP and 2 points in MS3). Interestingly, the segmentation using $U_{i}^{k^{\star}}$ does not seems to be improved a lot by the specific word bank (+1 point on S4 and MS3, and -5 points on ADE SP).
Appendix C TACO pseudo-code
The following pseudo-code details the exact computations performed by TACO. Note that this is the pseudo code for the case of a static frame, in the case of multiple frames (like a video) Eq 7 should be added to $\mathcal{L}$.
Algorithm 1 TACO algorithm
1:Input: Audio and image features ($X_{A}$, $X_{I}$), number of iterations $n_{\text{ite}}$, semantic anchors $\{b_{I}^{j}\}_{j=1}^{J}$, $\{b_{A}^{j}\}_{j=1}^{J}$, learning rate $\eta$, parameter $\beta_{p}$, FCCLIP model $s(.)$
2:Initialize: $U_{A}$, $U_{I}$, $V_{A}$, $V_{I}\sim\mathcal{N}(0,\Sigma)$
3:for iteration $t=1$ to $n_{\text{ite}}$ do
4:     Compute semantic components: $\mathcal{C}_{I}^{k}=\text{avg}(X_{I}\odot U_{I}^{k})$, $\mathcal{C}_{A}^{k}=\text{avg}(X_{A}\odot U_{A}^{k})$
5:     Compute semantic descriptors: $\mathcal{D}_{I^{k}}=\begin{pmatrix}\cos(\mathcal{C}_{I}^{k},b_{I}^{1})\\ \vdots\\ \cos(\mathcal{C}_{I}^{k},b_{I}^{J})\end{pmatrix}$, $\mathcal{D}_{A^{k}}=\begin{pmatrix}\cos(\mathcal{C}_{A}^{k},b_{A}^{1})\\ \vdots\\ \cos(\mathcal{C}_{A}^{k},b_{A}^{J})\end{pmatrix}$
6:     Compute total cost:
$\mathcal{L}=\|X_{A}-U_{A}V_{A}\|_{2}^{2}+\|X_{I}-U_{I}V_{I}\|_{2}^{2}+\beta_{p}\min_{k}\,\text{CE}(\mathcal{D}_{I^{k}},\mathcal{D}_{A^{k}})$
7:     Update parameters:
8:         $U_{A}\leftarrow U_{A}-\eta\frac{\partial\mathcal{L}}{\partial U_{A}}$
9:         $U_{I}\leftarrow U_{I}-\eta\frac{\partial\mathcal{L}}{\partial U_{I}}$
10:         $V_{A}\leftarrow V_{A}-\eta\frac{\partial\mathcal{L}}{\partial V_{A}}$
11:         $V_{I}\leftarrow V_{I}-\eta\frac{\partial\mathcal{L}}{\partial V_{I}}$
12:end for
13:$k^{\star}=\arg\min_{k}\text{CE}(\mathcal{D}_{I^{k}},\mathcal{D}_{A^{k}})$
14:segmentations $\leftarrow s(V_{I_{t}},X_{I})$
15:Return: segmentation[$k^{\star}$]
Figure 6: Illustration of the decomposition
The visualization of the decomposition is given in Figure 6. The input matrices are decomposed in 2 matrices: $U_{A}$ and $V_{A}$ and $U_{I}$ and $V_{I}$. The kth column of $U_{A}$ and $U_{I}$ that minimizes the cross entropy between the semantic descriptors is used to force the unify the 2 decompositions.
Appendix D Computational overhead
@@
Appendix E MACS analysis
The table 8 shows the cost (measured in MACs) of the image and audio encoder used in our method and concurrent ones. While our image encoder is larger, the efficiency of the CLAP audio encoder (HTSAT- Swin Transformer) significantly reduces the overall computation. As a result, our total encoding cost is lower than that of [18] and [34].
Method Image Encoder Audio Encoder Total
Hamilton et al. [18] $1.69\times 10^{10}$ $6.72\times 10^{10}$ $8.41\times 10^{10}$
Park et al. [34] $1.69\times 10^{10}$ $4.48\times 10^{10}$ $6.17\times 10^{10}$
Ours $3.44\times 10^{10}$ $6.31\times 10^{9}$ $4.07\times 10^{10}$
Table 8: Comparison of model parameter sizes for image and audio encoders.
Appendix F Metrics discussion
@@
Appendix G Semantic segmentation analysis
@@
Appendix H Qualitative results of single and multiple sources, sound-prompted segmentation
@@
Appendix I Typical failure cases
@@
Appendix J Dataset licence
The experiments were performed on datasets using different kinds of licenses: AVSBench (S4 and MS3) under Apache 2.0, ADESP and ADESP Semantic rely on ADE20k, which is under BSD 3-clause, and VGGSound, which is under Creative Commons Attribution 4.0, finally, the AVSS dataset is under an Apache 2.0 license.